How much heat is produced by the complete reaction of 4.47 kg

of nitromethane?


1
Expert's answer
2021-04-08T12:10:09-0400

The molar mass of nitromethane,"m= (12+3(1) +14+2(16)) g\/mol" =

61g/mol


Thus, the number of moles n in 5.56 kg of nitromethane is:

"n =4470\/61" =73.28moles


Nitromethane (CH3NO2) burns in air to produce significant amounts of heat.

2CH3NO2(l)+3/2O2(g)>2CO2(g)+3H2O(l)+N2(g) = -1418 kJ


According to the equation,

2 moles of nitromethane on combustion gives 1418 kJ of heat.

Thus, the heat produced by the combustion of 73.28 moles of nitromethane:


Heat produced="\frac{1418}{2} × 73.28 = 51955.52J"